OpenStudy (anonymous):

250 = 9600/(300-3x) solve for x

OpenStudy (anonymous):

x = 436/5

OpenStudy (saifoo.khan):

\[250(300-3x) = 9600\] 75000 - 750x = 9600 -750x = 9600 -75000 \[x = \frac{-436}{5}\]

OpenStudy (anonymous):

great thanks

OpenStudy (saifoo.khan):

welks! :)